A weather balloon is filled with helium to a volume of 270 L at 29 ∘C and 765 mmHg . The balloon ascends to an altitude where the pressure is 466 mmHg and the temperature is -42 ∘C. You may want to reference (Pages 233 - 234) Section 8.8 while completing this problem. Part A What is the volume of the balloon at this altitude? Express your answer using three significant figures.
Answer #1

Given:
Pi = 765 mmHg
Pf = 466 mmHg
Vi = 270 L
Ti = 29.0 oC
= (29.0+273) K
= 302 K
Tf = -42.0 oC
= (-42.0+273) K
= 231 K

use:
(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)
(765 mmHg*270 L)/(302.0 K) = (466 mmHg*Vf)/(231.0 K)
Vf = 339 L
Answer: 339 L